OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 4x − ___

OpenStudy (jdoe0001):

x intercepts at x = -6 and x = 2. so get the roots from there and expand \(\bf x = -6 \implies x+6=0 \qquad \textit{root }(x+6)\\ x = 2 \implies x-2 = 0 \qquad \textit{root }(x-2)\\ (x+6)(x-2) = x^2+4x + \square?\)

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

what is \(2\times 6\) ?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

thats it?